# Co-investment with Payoff Sharing Benefit Operators and Users in Network Design

### v1

@@ II-A Modeling the Mobility Network and Travel Demand
$+$ $(1-x_{e})\sum\limits_{a\in\mathcal{E}_{R}}\mathbb{1}_{\mathcal{E}^{a}_{e}}l_{a}(\frac{\gamma_{\mathrm{vot}}}{v_{A}}+\gamma_{\mathrm{A}}),\quad\forall m\in\mathcal{M},$ (2)
where $x_{e}$ is the service availability on edge $e$, $\gamma_{\mathrm{vot}}$ is the value of time, $\gamma_{\mathrm{a}}$ and $\gamma_{\mathrm{t}}$ are the distance-based price for using alternative modes and train service, respectively, $l_{e}$ is the travel distance on edge $e$, and $v_{a}$ and $v_{t}$ are the average speed of using alternative modes and the speed of trains, and $\mathbb{1}_{\mathcal{E}^{a}_{e}}$ denotes whether the edge $e$ is part of the alternative-mode edges that travelers will take if the train service on edge $a$ is not available. For the request $r_{m}$, a proportion $p_{m}\in[0,1]$ of trips will choose the train-prioritized route, which can be determined by:
@@ II-B1 Model structure and Variables
$\begin{cases}\min\left\{\sum\limits_{m\in\mathcal{M}}\mathbb{1}_{\mathcal{E}^{R}_{m}(e)}\alpha_{m}p_{m},c_{e}\right\},&\text{if }e\in\mathcal{E}_{R},\\ \sum\limits_{m\in\mathcal{M}}\mathbb{1}_{\mathcal{E}^{A}_{m}(e)}\alpha_{m}(1-p_{m})+\sum\limits_{a\in\mathcal{E}_{R}}\mathbb{1}_{\mathcal{E}^{a}_{e}}\Delta_{a},&\text{otherwise},\end{cases}$ (8)
where $p_{m}$ is the proportion of travelers with request $m$ who are willing to choose the train-prioritized routes. The boolean indicator functions $\mathbb{1}_{\mathcal{E}^{R}_{m}(e)}$ and $\mathbb{1}_{\mathcal{E}^{A}_{m}(e)}$ indicate whether edge $e$ is part of the edge set of the train-prioritized route and the alternative route, respectively. The amount of unserved travelers due to capacity constraints on the edge $a\in\mathcal{E}_{R}$ is defined as:
$\Delta_{a}:=\max\left\{0,\sum\limits_{m\in\mathcal{M}}\alpha_{m}p_{m}\mathbb{1}_{\mathcal{E}^{R}_{m}({a})}-c_{a}\right\}.$
@@ II-B3 Design Stage 1 - Non-Cooperative Network Design
The optimization problem (12) is a mixed integer nonlinear programming problem, with binary decision variables $(x^{t}_{e})_{e\in\mathcal{E}^{R}_{i}}$ and integer decision variables $(s^{t}_{e})_{e\in\mathcal{E}^{R}_{i}}$. All constraints are linearly related to the decision variables. For each design period $t$ and regional authority $i$, the optimization problem has $\mathcal{O}\left(E_{i}\right)$ decision variables and $\mathcal{O}\left(E_{i}+M_{i}\right)$ constraints, where $E_{i}$ denotes the number of rail edges in the region $i$. $M_{i}:=\sum_{r_{m}\in\mathcal{R}}\alpha_{m}\mathbb{1}\{o_{m}\in\mathcal{V}_{i}\vee d_{m}\in\mathcal{V}_{i}\}$ is the number of trips with the origin or the destination in the region $i$.
@@
$y_{e}=\sum_{m\in\mathcal{M}}\sum_{k\in\mathcal{K}}f_{m}^{k}\mathbb{1}_{\mathcal{E}^{k}_{m}(e)},$ (13d)

### v2

@@ II-A Modeling the Mobility Network and Travel Demand
$+$ $(1-x_{e})\sum\limits_{a\in\mathcal{E}_{R}}\mathds{1}_{\mathcal{E}^{a}_{e}}l_{a}(\frac{\gamma_{\mathrm{vot}}}{v_{A}}+\gamma_{\mathrm{A}}),\quad\forall m\in\mathcal{M},$ (2)
where $x_{e}$ is the service availability on edge $e$, $\gamma_{\mathrm{vot}}$ is the value of time, $\gamma_{\mathrm{a}}$ and $\gamma_{\mathrm{t}}$ are the distance-based price for using alternative modes and train service, respectively, $l_{e}$ is the travel distance on edge $e$, and $v_{a}$ and $v_{t}$ are the average speed of using alternative modes and the speed of trains, and $\mathds{1}_{\mathcal{E}^{a}_{e}}$ denotes whether the edge $e$ is part of the alternative-mode edges that travelers will take if the train service on edge $a$ is not available. For the request $r_{m}$, a proportion $p_{m}\in[0,1]$ of trips will choose the train-prioritized route, which can be determined by:
@@ II-B1 Model structure and Variables
$\begin{cases}\min\left\{\sum\limits_{m\in\mathcal{M}}\mathds{1}_{\mathcal{E}^{R}_{m}(e)}\alpha_{m}p_{m},c_{e}\right\},&\text{if }e\in\mathcal{E}_{R},\\ \sum\limits_{m\in\mathcal{M}}\mathds{1}_{\mathcal{E}^{A}_{m}(e)}\alpha_{m}(1-p_{m})+\sum\limits_{a\in\mathcal{E}_{R}}\mathds{1}_{\mathcal{E}^{a}_{e}}\Delta_{a},&\text{otherwise},\end{cases}$ (8)
where $p_{m}$ is the proportion of travelers with request $m$ who are willing to choose the train-prioritized routes. The boolean indicator functions $\mathds{1}_{\mathcal{E}^{R}_{m}(e)}$ and $\mathds{1}_{\mathcal{E}^{A}_{m}(e)}$ indicate whether edge $e$ is part of the edge set of the train-prioritized route and the alternative route, respectively. The amount of unserved travelers due to capacity constraints on the edge $a\in\mathcal{E}_{R}$ is defined as:
$\Delta_{a}:=\max\left\{0,\sum\limits_{m\in\mathcal{M}}\alpha_{m}p_{m}\mathds{1}_{\mathcal{E}^{R}_{m}({a})}-c_{a}\right\}.$
@@ II-B3 Design Stage 1 - Non-Cooperative Network Design
The optimization problem (12) is a mixed integer nonlinear programming problem, with binary decision variables $(x^{t}_{e})_{e\in\mathcal{E}^{R}_{i}}$ and integer decision variables $(s^{t}_{e})_{e\in\mathcal{E}^{R}_{i}}$. All constraints are linearly related to the decision variables. For each design period $t$ and regional authority $i$, the optimization problem has $\mathcal{O}\left(E_{i}\right)$ decision variables and $\mathcal{O}\left(E_{i}+M_{i}\right)$ constraints, where $E_{i}$ denotes the number of rail edges in the region $i$. $M_{i}:=\sum_{r_{m}\in\mathcal{R}}\alpha_{m}\mathds{1}\{o_{m}\in\mathcal{V}_{i}\vee d_{m}\in\mathcal{V}_{i}\}$ is the number of trips with the origin or the destination in the region $i$.
@@
$y_{e}=\sum_{m\in\mathcal{M}}\sum_{k\in\mathcal{K}}f_{m}^{k}\mathds{1}_{\mathcal{E}^{k}_{m}(e)},$ (13d)